# EditVerse: Unifying Image and Video Editing and Generation with In-Context Learning

### v2

@@ Abstract
Figure 1: The strong video editing performance of EditVerse emerges from a unified framework trained on a diverse set of mixed image and video data. This teaser visualizes a selection of supported image and video editing tasks (Full instructions are shown in the Appendix).
@@ 3.1 Interleaved Text and Vision Input
Following prior works (Kingma & Welling, 2013), we encode the RGB pixel-space videos and images into a learned spatio-temporally compressed latent space by training a convolutional Variational Autoencoder (VAE) capable of both feature extraction and reconstruction. Specifically, given an input image or video $I_{vision}$, the VAE compresses it into a continuous-valued latent representation with downsampling ratios $r_{T}$, $r_{H}$, $r_{W}$. Then, the vision features are patchified into a long token sequence with a $1\times 2\times 2$ kernel to get $X_{vision}\in\mathbb{R}^{L_{vision}\times C_{vision}}$ ($L_{vision}$ is the vision token number, $C_{text}$ is the channel dimension of vision feature). For a given text input $I_{text}$, we first generate text tokens using the Flan-T5-XXL (Chung et al., 2022) encoder. Then, we retain only the tokens that correspond directly to the input text, discarding the rest, yielding a final representation $X_{text}\in\mathbb{R}^{L_{text}\times C_{text}}$ ($L_{text}$ is the token count of $I_{text}$, $C_{text}$ is the channel dimension of Flan-T5-XXL), which saves computation while preserving the necessary information from text input.
@@ 3.2 Rotary Positional Embedding
To distinguish text, image, and video from each other and to indicate their sequential positions, we design a special Rotary Positional Embedding (RoPE) that incorporates spatial (height and width), sequential, and temporal dimensions (shown in Figure 2). For each of these four positional dimensions, we apply a separate RoPE computation. (1) Height and Width Dimensions: For images and video frames, the height and width dimensions correspond to the pixel coordinates, increasing incrementally from the top-left to the bottom-right corner (Polyak et al., 2025). The increment values reflect the number of pixels along the height and width axes. For text tokens, both dimensions are set to 0. (2) Sequential dimension: This dimension captures the global position within the overall sequence, starting from 0. The value is incremented by $1$ for each text token and image/video frame, up to the end of the sequence. (3) Temporal dimension: This dimension is used exclusively for video frames to encode their temporal order within a video clip. It begins at 0 and increases by $1$ for each subsequent frame. For text and image inputs, this dimension remains 0. The height, width, sequential, and temporal dimensions each compute a separate RoPE, which are assigned RoPE embedding dimensions of $56$, $56$, $12$, and $4$, respectively. To better support variable-length input, we use the NTK-aware interpolation (Peng et al., 2023) in RoPE calculation for context window extension.
@@ 3.3 Training and Inference Paradigm
$\mathcal{L}=\mathbb{E}_{t,X_{0},X_{1}}\left\|u_{\Theta}(X_{t},t)-(X_{1}-X_{0}))\right\|$
@@ 4 Data Pipeline
EditVerse is trained on large-scale data composed of: $1.9M$ image generation samples (around $2.0B$ tokens), $3.9M$ video generation samples (around $68.8B$ tokens), $6.0M$ image editing samples (around $12.6B$ tokens), and $288K$ video editing samples (around $10.2B$ tokens). Notably, the video editing datasets have significantly smaller sample number and are less diverse than image editing datasets. Thus, our architecture is specifically designed to transfer learned editing knowledge from the image domain to the video domain. We summarize the used datasets in Table 1 and provide further details below. A more detailed breakdown for each dataset are provided in Appendix Table A.4.
@@
Video Editing Data Pipeline. Due to the scarcity and inadequate quality of publicly available video editing datasets, we developed our pipeline to generate EditVerse Editing Data, which can be applied to obtain video editing pairs from any video input. (1) Object Removal and Addition. We first use Grounded-SAM-2 (Ravi et al., 2024; Ren et al., 2024) to extract object masks from the video. To improve the success rate of object removal, we filter candidates based on object name, total mask area, and detection confidence score. Next, we apply DiffuEraser (Li et al., 2025) to remove the masked objects. We use video pairs before and after the removal to construct object removal and addition data. (2) Object Replacement. Again, we obtain object masks from Grounded-SAM-2 (Ravi et al., 2024; Ren et al., 2024). Next, we leverage a Vision-Language Model (VLM) (Wang et al., 2024) to imagine plausible transformations of the object. We then use VACE (Jiang et al., 2025) to inpaint the masked region based on the VLM’s output. To improve the success rate, we apply dynamic adjustments to the mask’s shape and area, conditioned on the object’s size and geometry. (3) Style Transfer. Previous style transfer techniques primarily rely on inference-based video editing methods (Qi et al., 2023), which we found to be unreliable when handling more diverse styles (e.g., Minecraft style). To address this, we first apply an image style transfer model to edit the first frame, and then utilize VACE’s (Jiang et al., 2025) depth-guided first-frame-to-video feature to generate the full styled video. (4) Camera Change. We select $10$ camera movements and use ReCamMaster (Bai et al., 2025) to generate camera change data. (5) Mask Detection. We construct the mask detection dataset by converting object removal, object addition, and object replacement data using the prompt template: “I want to [edit prompt]. Detect the region that needs to be edited”. (6) Propagation. We build the propagation dataset by extracting the first edited frame from style transfer, object removal, object addition, and object replacement data.
@@ 5.3 Comparison to Previous Methods
Comparison on TGVE+. Following Movie Gen (Polyak et al., 2025), we evaluate EditVerse on TGVE+ (Singer et al., 2024b). Specifically, we follow previous works and measure (i) ViCLIPdir: text-video direction similarity, which evaluates the alignment between changes in captions and corresponding changes in the videos, and (ii) ViCLIPout: output similarity, which measures the similarity between the edited video and the output caption. As shown in Table 3, OursMethod surpasses previous methods on both metrics. It is worth noting that all TGVE+ videos are square, whereas our training data does not include any square video editing samples.
@@ 5.4 Analysis of Emergent Ability
We also find that EditVerse can surpass the ground-truth training data in both quality and success rate by leveraging knowledge from the image generation/editing and video generation domains. We show two examples for object removal and object replacement in Figure 7.
@@ A.2 Evaluation Details
User Study. To validate our automated metrics and directly measure human perceptual preferences, we conducted a comprehensive user study. We recruited detail-oriented participants to evaluate the performance of different AI video editing models. Using a web-based interface, participants were shown pairs of edited videos, labeled “Result 1” and “Result 2”, each generated by different models from the same source video and text instruction. Their task was to compare the two videos and choose among “Result 1 is better,” “Result 2 is better,” or “They are about the same” across three evaluation criteria: (1) Text-Instruction Alignment: Which video better follows the provided instruction? (2) Preservation of Unedited Regions: Are unmodified parts of the video accurately preserved, with minimal distortion or artifacts? Ideally, edits should only affect the intended object or region. Select the one that preserves better. (3) Aesthetic Quality: Which video is more visually appealing in terms of realism, smoothness, and overall perceptual quality? A video is considered the winner of a comparison if it achieves a majority of wins across these three criteria.
@@ A.4 Detailed Training Data
Unreal Engine videos with object-present and object-absent scenes.
@@
Internal depth-to-video dataset. Depth is detected with Depth Anything v2.
@@
Internal video-to-depth dataset. Depth is detected with Depth Anything v2.
@@
Internal sketch-to-video dataset. Sketch is detected with OpenCV Canny.
@@
Internal video-to-sketch dataset. Sketch is detected with OpenCV Canny.
@@
Internal first frame-to-video dataset.
@@
Internal pose-to-video dataset. Pose is detected with RTM-Pose.
@@
Internal video-to-pose dataset. Pose is detected with RTM-Pose.
@@
Internal high-quality identity-to-video dataset.
@@ A.5 Limitation and Future Work
Computational Cost. Our reliance on a full self-attention mechanism across a unified one-dimensional token sequence, while powerful for in-context learning, leads to significant computational overhead. The concatenation operation results in long sequence lengths, particularly for high-resolution or long-duration videos, which translates to high FLOPs and prolonged training and inference time. Future work could explore more efficient attention mechanisms to reduce the computational burden without compromising the model’s cross-modal learning capabilities.
@@
Dataset Quality. Although our data curation pipeline is crucial for enabling instruction-based video editing, the resulting dataset contains inherent noise. The editing instructions are often concise (averaging around 10 words) and may lack the detail required for highly complex or nuanced edits. Future efforts could focus on developing more advanced data generation and filtering techniques.
Generalist vs. Specialist Models. Our work highlights the potential of unified models, but it is plausible that for specific, well-defined tasks with abundant high-quality data (e.g., inpainting), a dedicated specialist model might still yield superior results. A systematic investigation into the trade-offs between our generalist framework and specialist models would be a fruitful direction for future research. This could help delineate the precise scenarios where a unified approach offers the most significant advantages and where specialized architectures remain preferable.

### v3

@@ Abstract
Figure 1: The strong video editing performance of EditVerse emerges from a unified framework trained on a diverse set of mixed image and video data. This teaser visualizes a selection of supported image and video editing tasks (Instructions in the Appendix). More results in our Project Page.
@@ 3.1 Interleaved Text and Vision Input
Following prior works (Kingma and Welling, 2013), we encode the RGB pixel-space videos and images into a learned spatio-temporally compressed latent space by training a convolutional Variational Autoencoder (VAE) capable of both feature extraction and reconstruction. Specifically, given an input image or video $I_{vision}$, the VAE compresses it into a continuous-valued latent representation with downsampling ratios $r_{T}$, $r_{H}$, $r_{W}$. Then, the vision features are patchified into a long token sequence with a $1\times 2\times 2$ kernel to get $X_{vision}\in\mathbb{R}^{L_{vision}\times C_{vision}}$ ($L_{vision}$ is the vision token number, $C_{vision}$ is the channel dimension of vision feature). For a given text input $I_{text}$, we first generate text tokens using the Flan-T5-XXL (Chung et al., 2022) encoder. Then, we retain only the tokens that correspond directly to the input text, discarding the rest, yielding a final representation $X_{text}\in\mathbb{R}^{L_{text}\times C_{text}}$ ($L_{text}$ is the token count of $I_{text}$, $C_{text}$ is the channel dimension of Flan-T5-XXL), which saves computation while preserving the necessary information from text input.
@@ 3.2 Rotary Positional Embedding
To distinguish text, image, and video from each other and to indicate their sequential positions, we design a special Rotary Positional Embedding (RoPE) that incorporates sequential, temporal and spatial (height and width) dimensions (shown in Figure 2). For each of these four positional dimensions, we apply a separate RoPE computation. (1) Sequential dimension: This dimension captures the global position within the overall sequence, starting from 0. The value is incremented by $1$ for each text token and image/video frame, up to the end of the sequence. (2) Temporal dimension: This dimension is used exclusively for video frames to encode their temporal order within a video clip. It begins at 0 and increases by $1$ for each subsequent frame. For text and image inputs, this dimension remains 0. (3) Height and Width Dimensions: For images and video frames, the height and width dimensions correspond to the pixel coordinates, increasing incrementally from the top-left to the bottom-right corner (Polyak et al., 2025). The increment values reflect the number of pixels along the height and width axes. For text tokens, both dimensions are set to 0. The sequential, temporal, height, and width dimensions each compute a separate RoPE, which are assigned RoPE embedding dimensions of $12$, 44, $56$, $56$ respectively. To better support variable-length input, we use the NTK-aware interpolation (Peng et al., 2023) in RoPE calculation for context window extension.
@@ 3.3 Training and Inference Paradigm
$\mathcal{L}=\mathbb{E}_{t,\mathbf{X}_{0},\mathbf{X}_{1}}\left|u_{\Theta}(\mathbf{X}_{t},t)-(\mathbf{X}_{1}-\mathbf{X}_{0})\right|^{2}$
@@ 4 Data Pipeline
EditVerse is trained on large-scale data composed of: $1.9M$ image generation samples (around $2.0B$ tokens), $3.9M$ video generation samples (around $68.8B$ tokens), $6.0M$ image editing samples (around $12.6B$ tokens), and $288K$ video editing samples (around $10.2B$ tokens). Notably, the video editing datasets have significantly smaller sample number and are less diverse than image editing datasets. Thus, our architecture is specifically designed to transfer learned editing knowledge from the image domain to the video domain. We summarize the used datasets in Table 1 and provide further details below. A more detailed breakdown for each dataset is provided in Appendix Table A.4.
@@
Video Editing Data Pipeline. Due to the scarcity and inadequate quality of publicly available video editing datasets, we developed our pipeline to generate EditVerse Editing Data, which can be applied to obtain video editing pairs from any video input. (1) Object Removal and Addition. We first use Grounded-SAM-2 (Ravi et al., 2024; Ren et al., 2024) to extract object masks from the video. To improve the success rate of object removal, we filter candidates based on object name, total mask area, and detection confidence score. Next, we apply DiffuEraser (Li et al., 2025) to remove the masked objects. We use video pairs before and after the removal to construct object removal and addition data. (2) Object Change. Again, we obtain object masks from Grounded-SAM-2 (Ravi et al., 2024; Ren et al., 2024). Next, we leverage a Vision-Language Model (VLM) (Wang et al., 2024) to imagine plausible transformations of the object. We then use VACE (Jiang et al., 2025) to inpaint the masked region based on the VLM’s output. To improve the success rate, we apply dynamic adjustments to the mask’s shape and area, conditioned on the object’s size and geometry. (3) Style Transfer. Previous style transfer techniques primarily rely on inference-based video editing methods (Qi et al., 2023), which we found to be unreliable when handling more diverse styles (e.g., Minecraft style). To address this, we first apply an image style transfer model to edit the first frame, and then utilize VACE’s (Jiang et al., 2025) depth-guided first-frame-to-video feature to generate the full styled video. (4) Camera Change. We select $10$ camera movements and use ReCamMaster (Bai et al., 2025a) to generate camera change data. (5) Mask Detection. We construct the mask detection dataset by converting object removal, object addition, and object change data using the prompt template: “I want to [edit prompt]. Detect the region that needs to be edited”. (6) Propagation. We build the propagation dataset by extracting the first edited frame from style transfer, object removal, object addition, and object change data.
@@ 5.3 Comparison to Previous Methods
Comparison on TGVE+. Following Movie Gen (Polyak et al., 2025), we evaluate EditVerse on TGVE+ (Singer et al., 2024b). Specifically, we follow previous works and measure (i) ViCLIPdir: text-video direction similarity, which evaluates the alignment between changes in captions and corresponding changes in the videos, and (ii) ViCLIPout: output similarity, which measures the similarity between the edited video and the output caption. As shown in Table 3, EditVerse surpasses previous methods on both metrics. It is worth noting that all TGVE+ videos are square, whereas our training data does not include any square video editing samples.
@@ 5.4 Analysis of Emergent Ability
We also find that EditVerse can surpass the ground-truth training data in both quality and success rate by leveraging knowledge from the image generation/editing and video generation domains. We show two examples for object removal and object change in Figure 7.
@@ 6 Conclusion
Acknowledgement. This project was supported in part by the Innovation and Technology Fund (UD-1a-MHP/213/24), Hong Kong S.A.R.
@@ A.2 Evaluation Details
User Study. To validate our automated metrics and directly measure human perceptual preferences, we conducted a comprehensive user study. The user study was outsourced to a professional external vendor, who recruited 20 annotators coming from diverse non-expert backgrounds. Each comparison pair was independently evaluated by 3 different annotators, and each annotator labeled 150 comparison pairs in total. To faithfully capture end-user preferences, we intentionally kept the instruction minimal and user-centric, providing only the brief prompt shown below, without additional technical guidance. This setup was chosen to reflect how typical users would judge the outputs rather than imposing task-specific expertise. Although we did not compute a formal inter-annotator agreement statistic, the redundancy of three independent judgments per pair helps mitigate noise and increases the robustness of the aggregated preferences. We recruited detail-oriented participants to evaluate the performance of different AI video editing models. Using a web-based interface, participants were shown pairs of edited videos, labeled “Result 1” and “Result 2”, each generated by different models from the same source video and text instruction. Their task was to compare the two videos and choose among “Result 1 is better,” “Result 2 is better,” or “They are about the same” across three evaluation criteria: (1) Text-Instruction Alignment: Which video better follows the provided instruction? (2) Preservation of Unedited Regions: Are unmodified parts of the video accurately preserved, with minimal distortion or artifacts? Ideally, edits should only affect the intended object or region. Select the one that preserves better. (3) Aesthetic Quality: Which video is more visually appealing in terms of realism, smoothness, and overall perceptual quality? A video is considered the winner of a comparison if it achieves a majority of wins across these three criteria. We find the user study shows a Pearson Correlation of 0.84 with automatic VLM evaluation, indicating a very strong positive correlation between the user study and VLM rankings.
@@ A.4 Detailed Training Data
Rendered videos with object-present and object-absent scenes.
@@
Depth-to-video dataset. Depth is detected with Depth Anything v2.
@@
Video-to-depth dataset. Depth is detected with Depth Anything v2.
@@
Sketch-to-video dataset. Sketch is detected with OpenCV Canny.
@@
Video-to-sketch dataset. Sketch is detected with OpenCV Canny.
@@
First frame-to-video dataset.
@@
Pose-to-video dataset. Pose is detected with RTM-Pose.
@@
Video-to-pose dataset. Pose is detected with RTM-Pose.
@@
High-quality identity-to-video dataset from OmniVCus (Cai et al., 2025b).
@@ A.5 Limitation and Future Work
Computational Cost. Our reliance on a full self-attention mechanism across a unified one-dimensional token sequence, while powerful for in-context learning, leads to significant computational overhead. The concatenation operation results in long sequence lengths, particularly for high-resolution or long-duration videos, which translates to high FLOPs and prolonged training and inference time. Future work could explore more efficient attention mechanisms to reduce the computational burden without compromising the model’s cross-modal learning capabilities. Our work was not specifically designed to improve efficiency, and the reported memory usage and inference time reflect the model’s raw, unoptimized performance. There are several practical ways to improve it, which we plan to explore in future work: (1) Using a higher-compression VAE. In our model, we use a VAE with relatively low compression ($8\times$ spatial downsampling), which leads to a large number of visual tokens. Recent VAEs can achieve $16\times$ spatial compression. If we replace our VAE with $16\times$ spatial compression rate models, the token length can be reduced to about one quarter, which directly lowers the attention cost. (2) Dynamic token selection. We can introduce a dynamic token selection mechanism that adaptively keeps only the most important context tokens and prunes redundant ones. This can reduce the effective sequence length for full attention, as explored in recent work FullDiT2 (He et al., 2025). (3) Distillation for faster inference. We can further apply model distillation and step distillation to reduce both the number of diffusion steps and the model size, which can noticeably speed up generation. (4) Future work could explore more efficient attention mechanisms (e.g., linear attention, Mamba attention) to reduce the computational burden without compromising the model’s cross-modal learning capabilities.
Figure 10: GPU memory usage and inference latency as functions of token length.
Token Length 17652 34342 63662 110822
Inference Time (second/20 steps) 21 48 129 349
Inference Time (second/50 steps) 52 118 323 873
GPU Memory (MB) 29511 32768 36663 44824
TFLOPs/second 67.9 58.2 39.4 25.4
Table 11: Inference efficiency across different token lengths
To assess the efficiency of our model with extended token length (as shown in Figure 10 and Table 11), we analyze the impact of token length on both efficiency and GPU memory usage. Experiment results demonstrate that resource consumption growth remains within a manageable and predictable range. Specifically, GPU Memory usage exhibits a strong linear relationship with token length. As the sequence length was scaled from 17,652 up to 110,822 tokens, the peak memory footprint increased modestly from 29.5 GB to 44.8 GB. This predictable and relatively slow scaling rate confirms that the extended context itself does not impose an unconstrained memory ceiling. Similarly, inference time shows a systematic, controllable increase with token length.
@@
Dataset Quality. Although our data curation pipeline is crucial for enabling instruction-based video editing, the resulting dataset contains inherent noise. The editing instructions are often concise (averaging around 10 words) and may lack the detail required for highly complex or nuanced edits. Furthermore, the automated methods used to generate editing pairs have an estimated success rate of around 65%, inevitably introducing imperfect or failed edits into the training corpus. Future efforts could focus on developing more advanced data generation and filtering techniques as mentioned in concurrent works (Wei et al., 2025; Mou et al., 2025; Bai et al., 2025b; Yang et al., 2025).
Generalist vs. Specialist Models. Our work highlights the potential of unified models, but it is plausible that for specific, well-defined tasks with abundant high-quality data (e.g., inpainting), a dedicated specialist model might still yield superior results.